Aug 28, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. Eastern

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Tentative Agenda

  1. Announcements
  2. Action Items
  3. FCAL HDFCAL log book
  4. BCAL HDBCAL log book
    1. SiPM mounting assignment for ch0-8 of downstream modules are connected to the incorrect bias voltage. This was noticed with Chris and Nick were assembling the final spare pair of readout modules.
    2. GlueX-doc-2540 Shayne Gryba
    3. GlueX-doc-2539 Laura Teigrob
  5. Reconstruction and Simulation
  6. Calibration
  7. Any other business


Attendees: Manuel, Elton, Will, Mark (Jlab); Zisis, Teagan (UofR); Christina (Athens); Matt (IU)

  1. Announcements
    1. ARR Review ongoing, closeout at 1 pm
      • No major issues raised so far
      • Some discussion of updating FCAL darkroom access with lock procedures according to the EH&S Manual 6111. Will wait for report and investigate options.
  2. Action Items
  3. FCAL HDFCAL log book
    1. Long data runs taken during week of solenoid testing
      • Aug 16-24, mostly continuous running, but runs interrupted by some crashes
      • Investigated data for abnormal channels
      • Replaced 13 pmts, with various problems. Most had HV drop to zero during the running. Some had high current, which overheated the bases.
      • Data recorded from the flashes, but also the HV, dynode voltage, temperature from bases for use in diagnostics.
      • The data were entered into the log and will be made available to Paul Smith to trouble shoot the failed bases
      • Matt: Need a global way to correlate DAQ and EPICS variables.
      • Mark: There has been some discussion on how to do this systematically, but depends on the number of variables involved in the study.
    2. Elton: Should investigate whether the inventory data base is a good match to tracking bases (talk to Sergey Pozdniakov)
    3. Elton: PR for VME module logic module has been placed, so that NIM electronics for the external cosmic-ray counters can be replaced with VME electronics and bring the setup to required standards.
    4. Bfield. No effect visible on the response of pmts due to the magnetic field during this time. No systematic study was performed yet. Note that the Fcal platform is pulled back from its running location.
  4. BCAL HDBCAL log book
    1. SiPM mounting assignment for ch0-8 of downstream modules are connected to the incorrect bias voltage. This was noticed with Chris and Nick were assembling the final spare pair of readout modules.
    2. Will: Data taken at 12 deg C, 23 deg C and will normalize at 18 deg C. For all three runs, we used the HV settings from the Apr 11 backup. Check response changes under the different temperature conditions.
    3. Cosmic-ray trigger
      • During solenoid test, the cosmic-ray trigger stopped producing triggers at about 600 A in the solenoid.
      • Interestingly, the bottom counters continued to count. 100-200 mV pulses were still produced and showed coincidences on the scope.
      • The top counter stopped counting and the pulses were reduced to < 10 mV. The counter was raised by 22" to move it away from the magnet. The HV was also increased from 1800 to 2100 V. The thresholds on the CFDs were reduced as much as possible, from about 70 mV down to about 60 mV. We took data with the counters in this configuration.
      • After the solenoid test, the HV was restored to 1800 V. The thresholds were left at the lower value.
      • Iron shields were ordered, but have not yet been installed, but should be in preparation for when the solenoid is reenergized.
    4. GlueX-doc-2540 Shayne Gryba
    5. GlueX-doc-2539 Laura Teigrob
    6. Shayne and Laura have placed their summer reports on the portal
      • Shayne's analysis seems to indicate that the asymmetry of north/south trigger rates is not due to the placement of the counters. Seems to point to a hardware trigger problem, that needs investigating.Simulation
  5. Simulation (Andrei)
    • Files are ready to start accumulating statistics with the updated geometry to match the installation
    • Elton: Consider raising the top cosmic-ray counter in simulation to match current setup. It was decided to do this and keep the top counter at its present location.
  6. Reconstruction (Tegan)
    • Installed GlueX software on CentOS
    • With new scripts will retry installation of scientific Linux
  7. Calibration (Zisis)
    • Meeting yesterday discussed plans for calibration of the many of the subdetectors. Plans for BCAL will be circulated for comments next week.
  8. Any other business